Eicosene is a molecular compound and nonelectrolyte with the empirical formula CH2. The freezing point of a solution prepared by dissolving 100mg of eicosene in 1.00g Benzene was 1.75 degrees C lower than the freezing point of pure benzene. What is the molar mass of eicosene if Kf for benzene is 4.90 Degrees C/m?
Answer #1

We know that ΔT f = iKf x m
Where

ΔT f = depression in freezing point

        = freezing point of pure solvent – freezing point of solution

        = 1.75 oC

K f = depression in freezing constant = 4.90 oC/m

i = vanthoff’s factor = 1 (it is nonelectrolyte)

m = molality of the solution

    = ( mass / Molar mass ) / weight of the solvent in Kg

   = ( 100x10-3 g / M) / (1.0x10-3g)

   = 100 / M

Plug the values we get

1.75 = 1x4.90x(100/M)

M = 280 g/mol

Therefore the molar mass of eicosene is 280 g/mol

Given empirical formula is CH2

Emperical formula mass = 12+(2x1) = 14

So n = molar mass / emperical mass

        = 280 / 14

       = 20

Therefore the molecular formula of the nonelectrolyte is 20(CH2) = C20H40